OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x2 = −4x − 5?

OpenStudy (anonymous):

Answer choices?

OpenStudy (anonymous):

(1, 0) and (−5, 0) (1, 5) (−1, 0) and (−5, 0) (−1, 0) and (5, 0)

OpenStudy (anonymous):

Okay well, To find the x intercepts, set y = 0, and solve. BTW, answer B cannot be correct because it doesn't have y = 0 So with y=0 you have ---> x2 - 4x - 5 = 0 This can be expressed as (x+1)*(x-5) = 0 This can be solved separately as x+1 = 0 and x-5 = 0 to find the two answers. So the answer would be D :)

OpenStudy (anonymous):

thanks ill put more up :)

OpenStudy (anonymous):

okay